Hello I would like to know a more in depth way of solving this problem
-3y= -2 - 7
-4x= -6y - 10.
I need to know how to solve it from beginning to end. please

Answer by checkley79(3341)

-3y=-2-7
-3y=-9
3y=9
y=9/3
y=3 ans.
Proof:
-3*3=-2-7
-9=-9
-4x=-6y=10
No unique solution because you have 2 unknowns & only 1 equation.